# Think Silently, Think Fast: Dynamic Latent Compression of LLM Reasoning Chains

### v4

@@ Abstract
Large Language Models (LLMs) achieve superior performance through Chain-of-Thought (CoT) reasoning, but these token-level reasoning chains are computationally expensive and inefficient. In this paper, we introduce Compressed Latent Reasoning (CoLaR), a novel framework that dynamically compresses reasoning processes in latent space through a two-stage training approach. First, during supervised fine-tuning, CoLaR extends beyond next-token prediction by incorporating an auxiliary next compressed embedding prediction objective. This process merges embeddings of consecutive tokens using a compression factor $c$ randomly sampled from a predefined range, and trains a specialized latent head to predict distributions of subsequent compressed embeddings. Second, we enhance CoLaR through reinforcement learning (RL) that leverages the latent head’s non-deterministic nature to explore diverse reasoning paths and exploit more compact ones. This approach enables CoLaR to: i) perform reasoning at a dense latent level (i.e., silently), substantially reducing reasoning chain length, and ii) dynamically adjust reasoning speed at inference time by simply prompting the desired compression factor. Extensive experiments across four mathematical reasoning datasets demonstrate that CoLaR achieves $14.1\%$ higher accuracy than latent-based baseline methods at comparable compression ratios, and reduces reasoning chain length by $53.3\%$ with only $4.8\%$ performance degradation compared to explicit CoT method. Moreover, when applied to more challenging mathematical reasoning tasks, our RL-enhanced CoLaR demonstrates performance gains of up to $5.4\%$ while dramatically reducing latent reasoning chain length by $82.8\%$. Project page: CoLaR-latent-reasoning.github.io.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in mathematical reasoning, particularly when employing Chain-of-Thought (CoT) prompting techniques Wei et al. (2022); Xiang et al. (2025). Recent advances have further highlighted the potential of this approach when combined with reinforcement learning on extended reasoning sequences Jaech et al. (2024); Guo et al. (2025); Team et al. (2025), revealing significant “aha-moments” in model performance. Despite these advances, a critical limitation persists: generating lengthy reasoning chains are computational costly, impeding efficiency and scalability. This inefficiency becomes particularly evident in real-world LLM applications, where extended reasoning chains create substantial server load, especially under high-concurrency conditions, underscoring the urgent need for more efficient reasoning methods.
@@
Several approaches have emerged to address this computational challenge Sui et al. (2025); Feng et al. (2025). One line of research focuses on enhancing efficiency at the token level, primarily by identifying and skipping less informative tokens Xia et al. (2025), prompting models to generate more concise reasoning steps Xu et al. (2025a); Aytes et al. (2025), or dynamically terminating reasoning when the model exhibits high confidence in a trial answer Yang et al. (2025). While valuable, these methods continue to operate on sparse token-based representations. A more promising direction explores reasoning within the dense latent space. Initial efforts attempt to “internalize” reasoning knowledge by curriculum learning Deng et al. (2024) or knowledge distillation Deng et al. (2023). Some works focus on the potential inside LLMs by looping or skipping some intermediate LLM layers Cheng and Van Durme (2024); Chen et al. (2025); Saunshi et al. (2025) to realize efficient reasoning. Recent innovations have introduced auto-regressive prediction of latent representations for efficient reasoning. Coconut Hao et al. (2024) proposes to gradually replace token-level reasoning with latent representations, while CODI Shen et al. (2025) employs self-distillation to transfer CoT knowledge into latent reasoning processes. However, these methods primarily utilize fixed-length reasoning chains, resulting in suboptimal efficiency and limited adaptability. Furthermore, to the best of our knowledge, all these latent-based methods employ deterministic latent reasoning processes, overlooking the potential benefits of exploration-exploitation capability may bring about.
@@
To overcome these limitations, we introduce Compressed Latent Reasoning (CoLaR), a novel framework that dynamically compresses LLM reasoning chains into latent space while preserving exploration-exploitation capabilities. Our approach utilizes an auxiliary next compressed embedding prediction task in supervised fine-tuning (SFT) stage. Specifically, at each training step, CoLaR first samples a random compression factor $c\in[1,c_{\max}]$ and merges the embeddings of $c$ consecutive reasoning tokens using our Embedding Compress module. A Latent Head is then trained to predict the next compressed embeddings from the LLM’s output hidden states, which is a fully parallelized process. During inference, CoLaR is capable to auto-regressively predict dense and informative latents with the Latent Head, and automatically determine when to terminate the reasoning process with LLM’s Language Head. Rather than predicting deterministic values, the Latent Head outputs a probability distribution that produces diverse reasoning pathways for a same question input. Based on this, we further enhance CoLaR through post-training with Group Relative Policy Optimization (GRPO) reinforcement learning algorithm Shao et al. (2024); Yu et al. (2025), which enables CoLaR to explore correct latent reasoning paths with diverse outputs and exploit those shorter ones.
@@
Our extensive evaluations on four grade-school level mathematical reasoning datasets (GSM8k Cobbe et al. (2021), GSM8k-hard Gao et al. (2023), SVAMP Patel et al. (2021), and MultiArith Roy and Roth (2015)) demonstrate that CoLaR achieves a $14.1\%\uparrow$ improvement in accuracy compared to state-of-the-art baseline methods at comparable compression ratios. Furthermore, CoLaR reduces reasoning chain length by $53.3\%\downarrow$ with only a $4.8\%\downarrow$ performance degradation relative to explicit CoT. Finally, experiments on a more challenging dataset MATH Hendrycks et al. (2021) demonstrates the potential of CoLaR to reinforcement learning, gaining up to $5.36\%\uparrow$ accuracy while reducing the length of reasoning chain significantly by $82.8\%\downarrow$.
@@ 2.1 Explicit LLM reasoning
Recent advances have demonstrated the strong reasoning capabilities of large language models (LLMs). The explicit reasoning approach, exemplified by Chain-of-Thought (CoT) reasoning Wei et al. (2022), propose to prompt LLMs to generate intermediate reasoning steps through sequential token prediction before generating answers Ye et al. (2025); Xiang et al. (2025); Tan et al. (2025); Zelikman et al. (2024). Subsequent work demonstrated that reinforcement learning techniques Shao et al. (2024); Yu et al. (2025); Yuan et al. (2025) can further improve performance on verifiable reasoning tasks like mathematical problem-solving, revealing an “aha-moment” that significantly boosts model’s performances with longer thinking process Jaech et al. (2024); Guo et al. (2025); Team et al. (2025).
However, the computational cost of processing these lengthy reasoning chains remains a significant bottleneck, motivating research into efficiency optimizations. Current solutions focus on identifying and skipping redundant tokens Xia et al. (2025); Yang et al. (2025) or encouraging more compact reasoning patterns using mathematical notations and programming language-like expressions Aytes et al. (2025); Xu et al. (2025a). While these methods reduce reasoning chain length, they are fundamentally limited by the sequential token prediction paradigm.
@@ 2.2 Latent LLM reasoning
The first direction, knowledge internalization, aims to embed reasoning capabilities directly into the model. iCoT-SI Deng et al. (2024) attempts to internalize reasoning knowledge by progressively removing explicit reasoning steps during training, while Pause Goyal et al. (2023) proposes training models to reason within specialized token embeddings.
The second direction exploits the hierarchical structure of transformer layers, with proposals to dynamically skip or repeat layer computations Saunshi et al. (2025); Chen et al. (2025); Cheng and Van Durme (2024); Su et al. (2025); Mohtashami et al. (2023). These methods aim to reduce computational cost by selectively processing different layers.
The third direction, and most relevant to our work, explores auto-regressive latent reasoning Xu et al. (2025b). Coconut Hao et al. (2024) pioneered this approach by replacing token sampling with hidden state concatenation for breadth-first reasoning, while CODI Shen et al. (2025) introduces an auto-regressive latent variable model through self-distillation. However, existing methods like Coconut and CODI are limited by their reliance on fixed-length reasoning chains due to the implicit nature of latent variables. Furthermore, they employ a deterministic approach to auto-regressive latent generation, neglecting the potential for exploration-exploitation strategies to further enhance model performance, particularly within a reinforcement learning framework.
@@ 3 Method
Figure 2: Our proposed method CoLaR consisting an LLM backbone and a Latent Head. During the SFT stage (left), for each training step, CoLaR first compresses embeddings $\mathbf{e}_{r}$ of the original reasoning chain into compressed embeddings $\mathbf{e}_{c}$ with a compression factor $c$ randomly selected from the range $[1,c_{max}]$. Then, CoLaR is trained to predict: i) the compressed reasoning embeddings via the Latent Head, and ii) the compressed reasoning tokens and answer tokens through the Language Head. During the RL stage (right), for every question input, CoLaR samples a group of $G$ outputs $o_{1:G}$ consisting of the latent reasoning chain and the predicted answer. We then calculate the relative rewards $a_{1:G}$ for each output, and the rewards are averaged on each token ($\bar{a}_{i}$), encouraging CoLaR to explore diverse latent reasoning pathways and exploit those more compact ones.
@@ 3.1 Reasoning token compression and understanding
As illustrated in Figure 2, the input to CoLaR can be represented as $\mathbf{e}=[\mathbf{e}_{q},\mathbf{e}_{c},\mathbf{e}_{a}]$, where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation. Here, $\mathbf{e}_{c}$ represents the compressed embeddings derived from $\mathbf{e}_{r}$, the embeddings of the original reasoning steps, and the length of compressed embeddings $L_{c}=\lceil\frac{L_{r}}{c}\rceil$. To achieve a dynamic test-time compression factor $c$, we begin each training step by randomly sampling $c\in[1,r_{\text{max}}]$. For every $r$ consecutive reasoning token embeddings $e_{r}^{k:k+r}$, the Embedding Compress module generates a compressed embedding $e_{c}^{k}$
A straightforward approach is to apply mean pooling directly to these embeddings. However, due to the high dimensionality of the embedding space (e.g., 2048 dimensions), embeddings from different tokens may be highly uncorrelated. Simply averaging these embeddings can distort the original distribution. For instance, consider two uncorrelated distributions $A\sim\mathcal{N}(\mu,\sigma^{2})$ and $B\sim\mathcal{N}(\mu,\sigma^{2})$; mean pooling would alter the original distribution to $\frac{A+B}{2}\sim\mathcal{N}(\mu,\frac{\sigma^{2}}{2})$, effectively scaling the variance by $\frac{1}{2}$. We found that, for most pre-trained LLMs, the distributions of embeddings are centered at $\mu\approx 0$. Thus, to prevent distortion of the original embedding distribution of LLMs, the Embedding Compress module only scales the sum of the $r$ embeddings by $\frac{1}{\sqrt{r}}$.
Intuitively, it could be difficult for LLMs to understand these compressed embeddings. A simple approach is to supervise $\mathcal{M}$ to predict answers with a language modeling loss, which enforces $\mathcal{M}$ to model answers with compressed embedding inputs. However, this objective provides supervision signals that are too sparse to converge to near-optimal performance. To address this issue, we train CoLaR to predict the compressed reasoning tokens. Ideally, when using a compression factor $r$, CoLaR should be able to read and predict tokens in groups of $r$. This means that for each compressed embedding input, CoLaR should be trained to predict all $r$ corresponding tokens. To approximate this multi-label classification task using the single-label prediction capability of LLM’s language model head, we randomly sample one token from each group of $r$ reasoning tokens $t_{r}^{k\times r:(k+1)\times r}$ as the ground-truth label. This approach trains the predicted logits to approximate a multimodal distribution that represents all potential tokens in each compressed group. This process could be formally represented as:
@@ 3.3 Exploration with reinforcement learning
With a trained Latent Head capable of exploration, we can sample diverse latent reasoning pathways and final answers for the same question $q$. We then apply Group Relative Policy Optimization (GRPO) algorithm Shao et al. (2024) to reinforce correct reasoning chains and answers while penalizing incorrect ones.
@@
We simply set $r_{i}$ to $1$ when an answer is correctly predicted and to 00 otherwise. Following DAPO Yu et al. (2025), we remove the KL-regularization term from original GRPO implementation for efficient training.
@@ 4.1 Experimental setup
Datasets and tasks. Our method is mainly trained and evaluated on GSM8k-Aug (Deng et al., 2023), an augmented version of the Grade-School level Math reasoning dataset GSM8k (Cobbe et al., 2021). GSM8k-Aug comprises approximately 385k training samples and 1k test samples. We also evaluate the trained methods on three out-of-domain math reasoning datasets: (1) GSM-Hard (Gao et al., 2023), a modified version of GSM8K with approximately 1k test samples featuring larger magnitude numbers, (2) SVAMP (Patel et al., 2021) and (3) MultiArith Roy and Roth (2015), two simpler math reasoning datasets with 1k and 600 test samples, respectively. Moreover, we train and evaluate our method on a more challenging dataset MATH Hendrycks et al. (2021), which consists of 7.5k training samples and 5k test samples, covering algebra, calculus, statistics, geometry, linear algebra and number theory. Following (Hao et al., 2024), we use two metrics: (1) Accuracy (Acc.), which measures the effectiveness of correctly predicting answers and (2) Reasoning chain length (# L), which measures efficiency by averaging the number of tokens/latents predicted in reasoning chains.
Baseline methods. We primarily compare against the following baselines: (1) CoT (Wei et al., 2022), which is fine-tuned on complete reasoning chains and answers, and performs token-level reasoning before predicting answers during inference; (2) iCoT Deng et al. (2024), which internalizes reasoning knowledge by gradually removing reasoning steps, and directly predicts answers during inference; (3) Coconut (Hao et al., 2024), which is fine-tuned with a curriculum process to gradually replace token-level reasoning steps with latent reasoning steps, and performs six steps of latent reasoning before predicting answers; and (4) Distill, is our reproduced version of CODI Shen et al. (2025) based on their implementation details as the code and model are not released. It self-distills token-level CoT into fixed-length latent reasoning steps, with an inference stage same to Coconut.
Implementation details. (1) Base model: unless otherwise specified, all experiments use a frozen Llama-3.2-1B-Instruct (Grattafiori et al., 2024) backbone with a trainable LoRA module (Hu et al., 2022). Following Coconut, all methods are initialized with weights from CoT-SFT to accelerate training. (2) Model checkpointing: for fair comparison, all models are trained for up to 50 epochs or 12 hours, whichever is reached first, and we choose the checkpoint that achieves the best accuracy on the validation set as the final model. (3) Hyper-parameters: we use the AdamW (Loshchilov and Hutter, 2017) optimizer with a fixed learning rate of 1e-4 and a weight decay of 1e-2 in SFT stage, and set the learning rate to 1e-6 in RL stage. We set $r_{max}=5$ to train CoLaR. During inference, we configure the LLM generation with a temperature of 1 and top-p of 0.9. All SFT training processes are conducted on a single machine with eight NVIDIA A100 GPUs and a total batch size of 256. For more implementation details, please refer to Appendix Section A.

### v5

@@ Abstract
Large Language Models (LLMs) achieve superior performance through Chain-of-Thought (CoT) reasoning, but these token-level reasoning chains are computationally expensive and inefficient. In this paper, we introduce Compressed Latent Reasoning (CoLaR), a novel framework that dynamically compresses reasoning processes in latent space through a two-stage training approach. First, during supervised fine-tuning, CoLaR extends beyond next-token prediction by incorporating an auxiliary next compressed embedding prediction objective. This process merges embeddings of consecutive tokens using a compression factor $c$ randomly sampled from a predefined range, and trains a specialized latent head to predict distributions of subsequent compressed embeddings. Second, we enhance CoLaR through reinforcement learning (RL) that leverages the latent head’s non-deterministic nature to explore diverse reasoning paths and exploit more compact ones. This approach enables CoLaR to: i) perform reasoning at a dense latent level (i.e., silently), substantially reducing reasoning chain length, and ii) dynamically adjust reasoning speed at inference time by simply prompting the desired compression factor. Extensive experiments across four mathematical reasoning datasets demonstrate that CoLaR achieves $14.1\%$ higher accuracy than latent-based baseline methods at comparable compression ratios, and reduces reasoning chain length by $53.3\%$ with only $4.8\%$ performance degradation compared to explicit CoT method. Moreover, when applied to more challenging mathematical reasoning tasks, our RL-enhanced CoLaR demonstrates performance gains of up to $5.4\%$ while dramatically reducing latent reasoning chain length by $82.8\%$.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in mathematical reasoning, particularly when employing Chain-of-Thought (CoT) prompting techniques cot ; metacot . Recent advances have further highlighted the potential of this approach when combined with reinforcement learning on extended reasoning sequences o1 ; r1 ; kimi15 , revealing significant “aha-moments” in model performance. Despite these advances, a critical limitation persists: generating lengthy reasoning chains are computational costly, impeding efficiency and scalability. This inefficiency becomes particularly evident in real-world LLM applications, where extended reasoning chains create substantial server load, especially under high-concurrency conditions, underscoring the urgent need for more efficient reasoning methods.
@@
Several approaches have emerged to address this computational challenge stopsurvey ; efficientsurvey . One line of research focuses on enhancing efficiency at the token level, primarily by identifying and skipping less informative tokens tokenskip , prompting models to generate more concise reasoning steps chainfodraft ; sketchofthought , or dynamically terminating reasoning when the model exhibits high confidence in a trial answer deer . While valuable, these methods continue to operate on sparse token-based representations. A more promising direction explores reasoning within the dense latent space. Initial efforts attempt to “internalize” reasoning knowledge by curriculum learning icot-si or knowledge distillation icot-kd . Some works focus on the potential inside LLMs by looping or skipping some intermediate LLM layers ccot ; innerthinking ; loop to realize efficient reasoning. Recent innovations have introduced auto-regressive prediction of latent representations for efficient reasoning. Coconut coconut proposes to gradually replace token-level reasoning with latent representations, while CODI codi employs self-distillation to transfer CoT knowledge into latent reasoning processes. However, these methods primarily utilize fixed-length reasoning chains, resulting in suboptimal efficiency and limited adaptability. Furthermore, to the best of our knowledge, all these latent-based methods employ deterministic latent reasoning processes, overlooking the potential benefits of exploration-exploitation capability may bring about.
@@
To overcome these limitations, we introduce Compressed Latent Reasoning (CoLaR), a novel framework that dynamically compresses LLM reasoning chains into latent space while preserving exploration-exploitation capabilities. Our approach utilizes an auxiliary next compressed embedding prediction task in supervised fine-tuning (SFT) stage. Specifically, at each training step, CoLaR first samples a random compression factor $c\in[1,c_{\max}]$ and merges the embeddings of $c$ consecutive reasoning tokens using our Embedding Compress module. A Latent Head is then trained to predict the next compressed embeddings from the LLM’s output hidden states, which is a fully parallelized process. During inference, CoLaR is capable to auto-regressively predict dense and informative latents with the Latent Head, and automatically determine when to terminate the reasoning process with LLM’s Language Head. Rather than predicting deterministic values, the Latent Head outputs a probability distribution that produces diverse reasoning pathways for a same question input. Based on this, we further enhance CoLaR through post-training with Group Relative Policy Optimization (GRPO) reinforcement learning algorithm grpo ; dapo , which enables CoLaR to explore correct latent reasoning paths with diverse outputs and exploit those shorter ones.
@@
Our extensive evaluations on four grade-school level mathematical reasoning datasets (GSM8k gsm8k , GSM8k-hard gsmhard , SVAMP svamp , and MultiArith multiarith ) demonstrate that CoLaR achieves a $14.1\%\uparrow$ improvement in accuracy compared to state-of-the-art baseline methods at comparable compression ratios. Furthermore, CoLaR reduces reasoning chain length by $53.3\%\downarrow$ with only a $4.8\%\downarrow$ performance degradation relative to explicit CoT. Finally, experiments on a more challenging dataset MATH math demonstrates the potential of CoLaR to reinforcement learning, gaining up to $5.36\%\uparrow$ accuracy while reducing the length of reasoning chain significantly by $82.8\%\downarrow$.
@@ 2.1 Explicit LLM reasoning
Recent advances have demonstrated the strong reasoning capabilities of large language models (LLMs). The explicit reasoning approach, exemplified by Chain-of-Thought (CoT) reasoning cot , propose to prompt LLMs to generate intermediate reasoning steps through sequential token prediction before generating answers limo ; metacot ; ttr ; quietstar . Subsequent work demonstrated that reinforcement learning techniques grpo ; dapo ; vapo can further improve performance on verifiable reasoning tasks like mathematical problem-solving, revealing an “aha-moment” that significantly boosts model’s performances with longer thinking process o1 ; r1 ; kimi15 .
However, the computational cost of processing these lengthy reasoning chains remains a significant bottleneck, motivating research into efficiency optimizations. Current solutions focus on identifying and skipping redundant tokens tokenskip ; deer or encouraging more compact reasoning patterns using mathematical notations and programming language-like expressions sketchofthought ; chainfodraft . While these methods reduce reasoning chain length, they are fundamentally limited by the sequential token prediction.
@@ 2.2 Latent LLM reasoning
The first direction, knowledge internalization, aims to embed reasoning capabilities directly into the model. iCoT-SI icot-si attempts to internalize reasoning knowledge by progressively removing explicit reasoning steps during training, while Pause pause proposes training models to reason within specialized token embeddings.
The second direction exploits the hierarchical structure of transformer layers, with proposals to dynamically skip or repeat layer computations loop ; innerthinking ; ccot ; tokenassorted ; cotformer . These methods aim to reduce computational cost by selectively processing different layers.
The third direction, and most relevant to our work, explores auto-regressive latent reasoning softcot . Coconut coconut pioneered this approach by replacing token sampling with hidden state concatenation for breadth-first reasoning, while CODI codi introduces an auto-regressive latent variable model through self-distillation. However, existing methods like Coconut and CODI are limited by their reliance on fixed-length reasoning chains due to the implicit nature of latent variables. Furthermore, they employ a deterministic approach to auto-regressive latent generation, neglecting the potential for exploration-exploitation strategies to further enhance model performance, particularly within a reinforcement learning framework.
@@ 3 Method
Figure 2: Our proposed method CoLaR consisting an LLM backbone and a Latent Head. During the SFT stage (left), for each training step, CoLaR first compresses embeddings $\mathbf{e}_{r}$ of the original reasoning chain into compressed embeddings $\mathbf{e}_{c}$ with a compression factor $c$ randomly selected from the range $1,c_{max}]$. Then, CoLaR is trained to predict: i) the compressed reasoning embeddings via the Latent Head, and ii) the compressed reasoning tokens and answer tokens through the Language Head. During the RL stage (right), for every question input, CoLaR samples a group of $G$ outputs $o_{1:G}$ consisting of the latent reasoning chain and the predicted answer. We then calculate the relative rewards $a_{1:G}$ for each output, and the rewards are averaged on each token ($\bar{a}_{i}$), encouraging CoLaR to explore diverse latent reasoning pathways and exploit those more compact ones.
@@ 3.1 Reasoning token compression and understanding
As illustrated in Figure 2, the input to CoLaR can be represented as $\mathbf{e}=[\mathbf{e}_{q},\mathbf{e}_{c},\mathbf{e}_{a}]$, where $\cdot$$\cdot$] denotes concatenation. Here, $\mathbf{e}_{c}$ represents the compressed embeddings derived from $\mathbf{e}_{r}$, the embeddings of the original reasoning steps, and the length of compressed embeddings $L_{c}=\lceil\frac{L_{r}}{c}\rceil$. To achieve a dynamic test-time compression factor $c$, we begin each training step by randomly sampling $c\in[1,r_{\text{max}}]$. For every $r$ consecutive reasoning token embeddings $e_{r}^{k:k+r}$, the Embedding Compress module generates a compressed embedding $e_{c}^{k}$ .
A straightforward approach is to apply mean pooling directly to these embeddings. However, due to the high dimensionality of the embedding space (e.g., 2048 dimensions), embeddings from different tokens may be highly uncorrelated. Simply averaging these embeddings can distort the original distribution. For instance, consider two uncorrelated distributions $A\sim\mathcal{N}(\mu,\sigma^{2})$ and $B\sim\mathcal{N}(\mu,\sigma^{2})$; mean pooling would alter the original distribution to $\frac{A+B}{2}\sim\mathcal{N}(\mu,\frac{\sigma^{2}}{2})$, effectively scaling the variance by $\frac{1}{2}$. We found that, for most pre-trained LLMs, the distributions of embeddings are centered at $\mu\approx 0$. Thus, to prevent distortion of the original embedding distribution of LLMs, the Embedding Compress module only scales the sum of the $c$ embeddings by $\frac{1}{\sqrt{c}}$.
Intuitively, it could be difficult for LLMs to understand these compressed embeddings. A simple approach is to supervise $\mathcal{M}$ to predict answers with a language modeling loss, which enforces $\mathcal{M}$ to model answers with compressed embedding inputs. However, this objective provides supervision signals that are too sparse to converge to near-optimal performance. To address this issue, we train CoLaR to predict the compressed reasoning tokens. Ideally, when using a compression factor $c$, CoLaR should be able to read and predict tokens in groups of $c$. This means that for each compressed embedding input, CoLaR should be trained to predict all $c$ corresponding tokens. To approximate this multi-label classification task using the single-label prediction capability of LLM’s language model head, we randomly sample one token from each group of $c$ reasoning tokens $t_{r}^{k\times c:(k+1)\times c}$ as the ground-truth label. This approach trains the predicted logits to approximate a multimodal distribution that represents all potential tokens in each compressed group. This process could be formally represented as:
@@ 3.3 Exploration with reinforcement learning
The next-compressed embedding prediction training enables latent reasoning chains to mimic original chain of thoughts. However, this paradigm inevitably limits the performance of latent reasoning models to their CoT teachers. To further explore the potential of CoLaR, we apply a reinforcement learning stage to our proposed method.
With the Latent Head trained to predict distributions, we can sample diverse latent reasoning pathways and final answers for the same question $q$. We then apply Group Relative Policy Optimization (GRPO) algorithm grpo to reinforce correct reasoning chains and answers while penalizing incorrect ones.
@@
We simply set $r_{i}$ to $1$ when an answer is correctly predicted and to 0 otherwise. Following DAPO dapo , we remove the KL-regularization term from original GRPO implementation for efficient training.
@@ 4.1 Experimental setup
Datasets and tasks. Our method is mainly trained and evaluated on GSM8k-Aug (icot-kd, ), an augmented version of the Grade-School level Math reasoning dataset GSM8k (gsm8k, ). GSM8k-Aug comprises approximately 385k training samples and 1k test samples. We also evaluate the trained methods on three out-of-domain math reasoning datasets: (1) GSM-Hard (gsmhard, ), a modified version of GSM8K with approximately 1k test samples featuring larger magnitude numbers, (2) SVAMP (svamp, ) and (3) MultiArith multiarith , two simpler math reasoning datasets with 1k and 600 test samples, respectively. Moreover, we train and evaluate our method on a more challenging dataset MATH math , which consists of 7.5k training samples and 5k test samples, covering algebra, calculus, statistics, geometry, linear algebra and number theory. Following (coconut, ), we use two metrics: (1) Accuracy (Acc.), which measures the effectiveness of correctly predicting answers and (2) Reasoning chain length (# L), which measures efficiency by averaging the number of tokens/latents predicted in reasoning chains.
Baseline methods. We primarily compare against the following baselines: (1) CoT (cot, ), which is fine-tuned on complete reasoning chains and answers, and performs token-level reasoning before predicting answers during inference; (2) iCoT icot-si , which internalizes reasoning knowledge by gradually removing reasoning steps, and directly predicts answers during inference; (3) Coconut (coconut, ), which is fine-tuned with a curriculum process to gradually replace token-level reasoning steps with latent reasoning steps, and performs six steps of latent reasoning before predicting answers; and (4) Distill, is our reproduced version of CODI codi based on their implementation details as the code and model are not released. It self-distills token-level CoT into fixed-length latent reasoning steps, with an inference stage same to Coconut.
Implementation details. (1) Base model: unless otherwise specified, all experiments use a frozen Llama-3.2-1B-Instruct (llama3, ) backbone with a trainable LoRA module (lora, ). Following Coconut, all methods are initialized with weights from CoT-SFT to accelerate training. (2) Model checkpointing: for fair comparison, all models are trained for up to 50 epochs or 12 hours, whichever is reached first, and we choose the checkpoint that achieves the best accuracy on the validation set as the final model. (3) Hyper-parameters: we use the AdamW (adamw, ) optimizer with a fixed learning rate of 1e-4 and a weight decay of 1e-2 in SFT stage, and set the learning rate to 1e-6 in RL stage. We set $r_{max}=5$ to train CoLaR. During inference, we configure the LLM generation with a temperature of 1 and top-p of 0.9. All SFT training processes are conducted a total batch size of 256. For more implementation details, please refer to Appendix Section A.